# PAC Apprenticeship Learning with Bayesian Active Inverse Reinforcement Learning

### v1

@@ Appendix B Theoretical Analysis
$>\min\left\{\frac{1}{|\mathcal{A}|},\frac{\delta}{(|\mathcal{A}|-1)|\mathcal{S}|}\right\}\frac{1-e^{-\beta(1-\gamma)\epsilon}}{4|\mathcal{A}|^{2}(|\mathcal{A}|-1)^{2}}$
$=\frac{1-e^{-\beta(1-\gamma)\epsilon}}{4|\mathcal{A}|^{2}(|\mathcal{A}|-1)^{3}|\mathcal{S}|}.$
The first inequality follows from the fact that all three terms in the sum are non-negative. The second inequality just plugs in results previously derived in this proof. The final step resolves the minimum as its second term using the assumption that $\delta\leq\frac{1}{2}$ and $|\mathcal{A}|\geq 2$.
Since the random variable $E_{s,a^{\prime}}$ is coarser than the variable $E^{\pi^{\text{A}}}$, which is the collection of the variables $E_{s,a}$ across all state-action pairs, we have $I(A;E)\geq I(A;E_{s,a^{\prime}})$, which completes the proof. ∎
@@
Let $H_{n}$ denote the expected entropy of $E$ after $n$ steps of active learning. By the properties of entropy and information gain:
@@
$=h_{\text{prior}}-N\cdot{\text{EIG}}_{\text{min}}(\epsilon,\delta)$
@@
$N\leq\frac{h_{\text{prior}}}{{\text{EIG}}_{\text{min}}(\epsilon,\delta)}$ (25)
The result follows by substituting the expression for ${\text{EIG}}_{\text{min}}(\epsilon,\delta)$ from Theorem 1. ∎
@@
$\log(3)|\mathcal{S}||\mathcal{A}|^{2}/{\text{EIG}}_{\text{min}}(\epsilon,\delta)=\frac{4\log(3)|\mathcal{A}|^{4}(|\mathcal{A}|-1)^{3}(|\mathcal{S}|^{2}}{\delta e^{-\beta(1-\gamma)\epsilon}(1-e^{-\beta(1-\gamma)\epsilon/2})^{2}}.$ (26)
@@
The random variable $E=\{E^{\pi^{\text{A}}}\;\forall{\pi^{\text{A}}}\}$ aggregates $|\mathcal{S}||\mathcal{A}|^{2}$ ternary random variables (in each state, apprentice policies can take $|\mathcal{A}|$ actions, and we are considering each action’s immediate regret relative to $|\mathcal{A}|-1<|\mathcal{A}|$ alternative actions), so $E$ can take at most $3^{|\mathcal{S}||\mathcal{A}|}$ values. Thus, its maximum entropy is $\log(1/3^{|\mathcal{S}||\mathcal{A}|^{2}})=|\mathcal{S}||\mathcal{A}|^{2}\log(3)$. The result follows by plugging this maximum entropy into Theorem 2. ∎
@@ B.1 Notes on possible improvements
B.1.1 Tighter bound for large state spaces
Note that the bound from Lemma 2 can be tightened if the state space is large and only a subset is reachable within an effective horizon. In that case $|S|$ can be replaced by the number of states reachable from the initial states within $1/(1-\gamma)$ steps. Also, if there is a limited horizon, or the apprentice policy always reaches a terminal state in a certain number of steps, we can also have fewer steps than that (and can further reduce the $(1-\gamma)$ factor in the exponent of our theoretical results.
B.1.2 Static policy
The bound includes the entropy of each action in each state. In fact, it may be enough to focus on a single action in each state, since we want to identify only a particular PAC policy, rather than reducing entropy of all of the components of $E(s)$ in every state. This should allow us to exclude a factor of $|A|$ from the bound on the expected number of steps to reach the PAC condition.
Appendix C Accounting for visitation frequencies
A practically useful acquisition function should account for one more important point: in order to get a policy with low expected regret, we do not need to reduce the expected immediate regret of all points, but just those that are likely to get visited by the apprentice policy $\pi^{\text{A}}$}. Let $\nu_{A}(s):=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}[s_{t}=s]]$ be the discounted expected occupancy of state $s$. Then, we can set
$\tilde{\alpha}(s):=\nu_{A}(s)I(A_{s},E_{s}).$ (27)
to be the acquisition function for querying single states.
However, we also wish to have an acquisition function for collecting full trajectories. A naive approach employed by some prior work (Kweon et al., 2023) would be summing the individual
$\alpha_{n}^{\prime}(s_{0}):=\mathbb{E}_{r|\mathcal{D}_{n}}\biggl{[}\mathbb{E}_{\tau|r}\bigl{[}\sum_{s\in\tau}\tilde{\alpha}(s)\bigr{]}\biggr{]}$ (28)
However, the sum in Eq. (28) neglects correlation between the regrets in different states (and, worse, autocorrelation if a state is visited multiple times). We can instead estimate the full expected information gain about our variable $E$ from the new expert trajectory. Then, dropping the weighing for the moment,
$\text{EIG}_{E}(s_{0})=\mathbb{E}_{\tau,E}\bigl{[}\log p(\tau|E)-\log p(\tau)\bigr{]}.$ (29)
We assume that the Bayesian IRL method we use to estimate this is able to give us samples from the current posterior over Q-values. Given a Q-value and an initial state, we can sample the corresponding hypothetical expert trajectories $\tau|Q$. Also, $E$ is a cheaply-computable function of Q, so we can also easily convert the samples of Q into samples of $E$. Then, the only remaining challenge in computing $\text{EIG}_{E}$ is estimating $p(\tau|E)$.
If the state space is small and we have a lot of Q samples, we can estimate $p(\tau|E)=\frac{1}{|\mathcal{Q}_{E}|}\sum_{Q\in\mathcal{Q}_{E}}p(\tau|Q)$. However, note that for a given policy, there are $3^{|}S|$ possible values of $E^{\pi}$, so even for a moderate size of the state space, the number of Q corresponding to each $E$ could be small. At the same time, the components of $E^{\pi}$ corresponding to states far away from the trajectory are unlikely to share much mutual information with the trajectory. Thus we suggest using the bound
$I(\tau,E)\geq I(\tau,E^{\pi}_{\tau})$ (30)
where $E^{\pi}_{\tau}:=\bigl{(}E_{s}\bigr{)}_{s\in\mathcal{S}_{\tau}}$ for $\mathcal{S}_{\tau}$ being some neighbourhood of $\tau$ in the state space, including all states on the trajectory $\tau$ plus states that can be quickly reached from the trajectory.
To again add the weighing, we can note that since the transition probabilities are the same for both components,
$\log p(\tau|E)-\log p(\tau)=\sum_{s,a\in\tau}\log p(a|s;E)-\log p(\tau|s)$ (31)
which naturally allows us to re-introduce the weights as
$\sum_{s,a\in\tau}\nu_{\pi}(s,a)(\log p(a|s;E)-\log p(\tau|s))$ (32)
resulting in an acquisition function
$\text{EIG}_{\nu\text{-PAC}}(s_{0})=\mathbb{E}_{\tau,E}\bigl{[}\sum_{s,a\in\tau}\nu_{\pi}(s,a)(\log p(a|s;E_{\tau})-\log p(\tau|s))\bigr{]}.$ (33)
Appendix D Alternative Acquisition Functions
In this appendix, we discuss alternative formulations of acquisition functions for active IRL, which could help to see the reasoning process that led to PACEIGPAC-EIG, and explain why some, possibly more, obvious alternatives were not chosen.
D.1 From Apprentice Return to Regret
@@
D.2 The Challenge of Direct Regret Optimization
@@
D.3 Information Gain About Regret
@@ D.5 Discretization and Connection to PAC-EIG
    Two-bucket PAC discretization: Using just two buckets – acceptable regret $0,\epsilon]$ and unacceptable regret $(\epsilon,\infty)$ – directly captures what matters for PAC guarantees. However, the theoretical arguments as shown above do not directly apply to this case, since we loose the middle bucket that ensures separation between the two sufficiently different expert action distributions. However, we do think a weaker result could be proven even for this case.
The PAC discretization is not only computationally more tractable but also theoretically motivated: it focuses information gathering on exactly what we need to know to provide formal reliability guarantees.
D.6 Summary
@@
Appendix E Experiment details
E.1 Basic parameter values
@@ E.4 Metrics
E.5 Implementation
The experiments were implemented using Python 3.10, PyTorch 2.5.1, and Pyro 1.8.6. We will publish our full code for both the experiments and the associated result analysis on Github once the anonymity requirement is lifted.
E.6 Timing
@@
Appendix F Notation overview

### v2

@@ Appendix B Theoretical Analysis
$>\min\left\{\frac{1}{|\mathcal{A}|},\frac{\delta}{(|\mathcal{A}|-1)|\mathcal{S}|}\right\}\frac{(1-e^{-\beta(1-\gamma)\epsilon})^{2}}{4|\mathcal{A}|^{2}(|\mathcal{A}|-1)^{2}}$
$=\frac{\delta(1-e^{-\beta(1-\gamma)\epsilon})^{2}}{4|\mathcal{A}|^{2}(|\mathcal{A}|-1)^{3}|\mathcal{S}|}.$
The first inequality follows from dropping the non-negative term corresponding to the event $e_{\text{a}}$ and taking the minimum of the remaining two probability terms. The second inequality just plugs in results previously derived in this proof. The final step resolves the minimum as its second term using the assumption that $\delta\leq\frac{1}{2}$ and $|\mathcal{A}|\geq 2$.
Since the random variable $E^{\pi^{\text{A}}}_{s,a^{\prime}}$ is coarser than the variable $E^{\pi^{\text{A}}}$, which is the collection of the variables $E^{\pi^{\text{A}}}_{s,a}$ across all state-action pairs, we have $I(A;E^{\pi^{\text{A}}})\geq I(A;E^{\pi^{\text{A}}}_{s,a^{\prime}})$, which completes the proof. ∎
@@
Let $H_{n}$ denote the expected entropy of $E$ after $n$ steps of active learning; in particular, $H_{0}$ is the entropy of the prior distribution over $E$. By the properties of entropy and information gain:
@@
$N\leq\frac{H_{0}}{{\text{EIG}}_{\text{min}}(\epsilon,\delta)}$ (24)
The result follows by substituting the expression for ${\text{EIG}}_{\text{min}}(\epsilon,\delta)$ from Theorem 1 and the fact that $H_{0}\leq h_{\text{max}}$ since $h_{\text{max}}$ is an upper bound on the prior entropy of $E$. ∎
@@
$\log(3)|\mathcal{S}||\mathcal{A}|^{2}/{\text{EIG}}_{\text{min}}(\epsilon,\delta)=\frac{4\log(3)|\mathcal{A}|^{3}(|\mathcal{A}|-1)^{4}|\mathcal{S}|^{2}}{\delta(1-e^{-\beta(1-\gamma)\epsilon})^{2}}.$ (25)
@@
The random variable $E=\{E^{\pi^{\text{A}}}\;\forall{\pi^{\text{A}}}\}$ aggregates $|\mathcal{S}||\mathcal{A}|(|\mathcal{A}|-1)$ ternary random variables (in each state, apprentice policies can take $|\mathcal{A}|$ actions, and we are considering each action’s immediate regret relative to $|\mathcal{A}|-1$ alternative actions), so $E$ can take at most $3^{|\mathcal{S}||\mathcal{A}|(|\mathcal{A}|-1)}$ values. Thus, its maximum entropy is $\log(1/3^{|\mathcal{S}||\mathcal{A}|(|\mathcal{A}|-1)})=|\mathcal{S}||\mathcal{A}|(|\mathcal{A}|-1)\log(3)$. The result follows by plugging this maximum entropy into Theorem 2. ∎
@@ B.1 Notes on possible improvements
B.1.1 Tighter bound for large state spaces and limited horizon
Note that the bound from Lemma 2 can be tightened if the state space is large and only a subset is reachable within an effective horizon. In that case $|S|$ can be replaced by the number of states reachable from the initial states within $1/(1-\gamma)$ steps. Also, if there is a limited time horizon, or the apprentice policy always reaches a terminal state in a certain number of steps, we can use this to tighten the $1/(1-\gamma)$ effective horizon and the associated $(1-\gamma)$ factor in the exponent of our theoretical results.
B.1.2 Dependence on the number of actions
The bound also considers reducing the entropy for every possible apprentice policy. If aiming for a tight PAC bound, the training will generally be able to quickly reduce the space of plausible apprentice policies and then just keep refining the bound for that policy in most states reducing the dependence on the number of actions in practice.
Appendix C Regret-based Acquisition Functions: Journey toward PAC-EIG
In this appendix, we discuss alternative formulations of acquisition functions for regret-focused active IRL, which could help to see the reasoning process that led to PAC-EIG, and explain why some seemingly obvious alternatives were not chosen.
C.1 From Apprentice Return to Regret
@@
C.2 The Challenge of Direct Regret Optimization
@@
C.3 Information Gain About Regret
@@ C.5 Discretization and Connection to PAC-EIG
    Two-bucket PAC discretization: Using just two buckets – acceptable regret $0,\epsilon]$ and unacceptable regret $(\epsilon,\infty)$ – directly captures what matters for PAC guarantees. However, the theoretical arguments as shown above do not directly apply to this case, since we loose the middle bucket that ensures separation between the two sufficiently different expert action distributions.
Our three-bucket PAC discretization is not only computationally more tractable but also theoretically motivated: it focuses information gathering on exactly what we need to know to provide formal reliability guarantees.
C.6 Summary
@@
Appendix D Experiment details
D.1 Basic parameter values
@@ D.4 Metrics
D.5 Implementation
The experiments were implemented using Python 3.10, PyTorch 2.5.1, and Pyro 1.8.6. We will publish our full code for both the experiments and the associated result analysis on Github.
D.6 Timing
@@
Appendix E Notation overview